What is the molar mass of a gas which has a mass of 834 mg, a volume of 267 mg, a temperature of 85°C, and a pressure of 995 mm Hg?
Answer #1

Given:

P = 995.0 mm Hg

= (995.0/760) atm

= 1.3092 atm

V = 267.0 mL

= (267.0/1000) L

= 0.267 L

T = 85.0 oC

= (85.0+273) K

= 358 K

find number of moles using:

P * V = n*R*T

1.3092 atm * 0.267 L = n * 0.08206 atm.L/mol.K * 358 K

n = 1.19*10^-2 mol

mass(solute)= 834 mg

= 0.834 g

use:

number of mol = mass / molar mass

1.19*10^-2 mol = (0.834 g)/molar mass

molar mass = 70.09 g/mol

Answer: 70.09 g/mol